if the median of the data is 63, find value of x.
29, 32, 48, 50, x, x+2, 72, 78, 84, 95.

Answer by mathmate(429)


Question:
if the median of the data is 63, find value of x.
29, 32, 48, 50, x, x+2, 72, 78, 84, 95.

Solution:

By understanding the definition of median, we would be able to find the value of x.

Median is the middle number, i.e. there should be as much data less than the median as there is above.

Since the median is 63, and there are 10 numbers, then 63, by definition of the median, is the average (mean) between the two middle numbers. Since the given numbers are sorted in ascending order, we will assume that x and x+2 are both sorted.

Thus the average of the two middle numbers is (x+(x+2))/2=x+1 = median.
Since we are given the median as 63, we can solve for x
(x+1)=63 => x=62.
Thus the series of numbers is:
29, 32, 48, 50, 62, 64, 72, 78, 84, 95.